OpenStudy (anonymous):

Find the vertical and horizontal intercepts of the function f(x) = (x-1)(2x+1)

OpenStudy (anonymous):

I got (-1, 0) for the vertical intercept but when I graph it on the calculator, the y-intercept is 1 y = (0-1)(2*0+1) (-1)(1)=-1

OpenStudy (anonymous):

The y-intercept is 0 on the calculator

OpenStudy (ranga):

y-intercept is when x = 0: y = (0-1)(0+1) = -1. (0,-1). x-intercept is when y = 0: 0 = (x-1)(2x+1). x = 1 or -0.5. (-0.5,0) & (1,0).

OpenStudy (anonymous):

Why would I get 0 on the calculator for the y-intercept?

OpenStudy (anonymous):

I used 2nd and trace with the zero function

OpenStudy (ranga):

I don't have a calculator with that functionality and so I don't know what you entered wrong. But there must be some mistake. Perhaps did not make use of parenthesis correctly or something like that. But the above answers are correct.

OpenStudy (anonymous):

Where does it show the intercepts on wolfram alpha?

OpenStudy (ranga):

Look at the first graph. It cuts the y-axis at -1 (vertical intercept) and the x axis at -0.5 and 1.0.

OpenStudy (anonymous):

Yeah, I see it